Linux VAR HOWTO

Martin Michlmayr, tbm@cyrius.com, http://www.cyrius.com/tbm v1.03, 11 April 1997 This document contains a listing of service companies which do not manufacture hardware or create packaged software, but add value to existing products. It is maintained by Martin Michlmayr <tbm@cyrius.com>.

Contents

1 About this document 1

2 Copyright Information 2

3 Preconfigured hardware 2 3.1 ASL Workstations, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.2 Apache Digital Corporation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.3 bee Baastrup EDV-Entwicklung GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.4 Centurion Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.5 Hard Data Ltd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.6 International Technology Sales . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.7 Linux Hardware Solutions, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.8 Nene Information Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.9 Netactive Systems, Ltd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.10 Net Express . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.11 Pacific Internet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.12 PRONETICS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.13 Sound Software Ltd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.14 SW Technology . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.15 Telenet System Solutions, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

1 About this document

This is the Linux VAR HOWTO. It contains a listing of service companies which do not manufacture hardware or create packaged software, but add value to existing products. This added value can take many forms, such as installation of Linux on computer hardware, configuration, or custom applications. It is maintained by Martin Michlmayr <tbm@cyrius.com>.

If you need to know more about the Linux Documentation Project or about Linux HOWTO’s, feel free to contact the supervisor Greg Hankins <gregh@sunsite.unc.edu>.

Greg Hankins will post the listing to several national and international newsgroups on a monthly basis. In addition, the Linux VAR HOWTO can be found on the World Wide Web at http://www.cyrius.com/tbm/VAR-HOWTO. New versions of the Linux VAR HOWTO are always placed at this site first, so please be sure to check if the copy you are reading is still up to date!

2 Copyright Information

This HOWTO is Copyright \begin{matrix}c \\ ◯ \end{matrix} 1997 by Martin Michlmayr.

A verbatim copy may be reproduced or distributed in any medium physical or electronic without permission of the author. Translations are similarly permitted without express permission if it includes a notice on who translated it.

Short quotes may be used without prior consent by the author. Derivative work and partial distributions of the Linux VAR HOWTO must be accompanied with either a verbatim copy of this file or a pointer to the verbatim copy.

Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions.

In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs.

We further want that all information provided in the HOWTOs is disseminated. If you have questions, please contact Greg Hankins, the Linux HOWTO coordinator, at gregh@sunsite.unc.edu.

3 Preconfigured hardware

This section contains companies selling preconfigured hardware for the purpose of running Linux.

Companies selling such hardware are invited to fill out the following form and contact me at tbm@cyrius.com.

     Name:
     Address:
     Phone:
     Fax:
     EMail:
     URL:
     Contact (Name of contact person):
     Architecture (Intel, Alpha, SPARC):
     Linux distribution (e.g. Debian GNU/Linux, Red Hat Linux):
     Special expertise:

3.1 ASL Workstations, Inc.

          2342 Shade Tree Lane
          San Jose, CA 95131
          U.S.A.

Phone:

+1 408 923 3837 Fax:

+1 408 923 4615 EMail:

info@aslab.com URL:

http://www.aslab.com Contact:

Jeff Nguyen Architecture:

Intel and Alpha. Linux distribution:

Craftworks Linux and Red Hat Linux. Special expertise:

Custom built high end servers and workstations with up to 4 GB of memory and support for RAID level 0, 1 and 5. Custom built rackmount enclosures are available too.

Last modified:

March 6, 1997.

3.2 Apache Digital Corporation

          777 Main Ave. Suite 211
          Durango, CO 81301
          U.S.A.

Phone:

+1 970 259 8153 Fax:

+1 970 259 8145 EMail:

info@apache.com URL:

http://www.apache.com Contact:

Mark Wauchope Architecture:

Intel, Alpha and SPARC. Linux distribution:

Caldera OpenLinux, Red Hat Linux and Slackware. Special expertise:

Custom service configuration available including innd, httpd, ftpd, telnetd, sendmail, POP 3, radius, procmail, etc. Especially useful to ISPs and network administrators.

Last modified:

March 7, 1997.

3.3 bee Baastrup EDV-Entwicklung GmbH

          Mallinkrodtstra{\ss}e 7
          D-44145 Dortmund
          Germany

Phone:

+49 231 9812060 Fax:

+49 231 9812062 EMail:

info@bee.de URL:

http://www.bee.de Contact:

Jörn Baastrup, Jürgen Stanke, Wiebke Schrader. Architecture:

Intel.

Linux distribution:

• Germany: DLD and S.u.S.E.

• International: Red Hat Linux

Special expertise:

• Configurated Linux servers and workstations

• bee-iX11-terminals (diskless Linux workstations)

• Rack-mountable Linux workstations for scientific needs

• bee-Linux-DialIn-Servers and Routers, based in Cyclades cards and ICN ISDN cards

• Distribution of different hardware components especially for Linux users: Cyclades serial multiport cards, ICN ISDN cards, DPT SCSI host adapters and RAID controler, Berkshire hardware watchdog cards, ELSA graphic adapters for the XFree86 server and/or Xinside X11 server, commercial software products for Linux.

Last modified:

March 6, 1997.

3.4 Centurion Services

          9351 Blundell Road
          Richmond, BC V6Y 1K5
          Canada

Phone:

+1 604 279 1857 Fax:

+1 604 279 1800 EMail:

censvc@ibm.net URL:

http://www.ipipeline.net/centurion Contact:

Jan Walter Architecture:

Intel and Alpha. Linux distribution:

Caldera OpenLinux, Red Hat Linux and Slackware.

Special expertise:

Integration with non-Unix systems. Intranet and Internet databases and mail. Last modified:

April 2, 1997.

3.5 Hard Data Ltd.

          11060 - 166 Avenue
          Edmonton, AB T5X 1Y3
          Canada

Phone:

+1 403 456 1510 Fax:

+1 403 456 1510 EMail:

harddata@bigfoot.com Contact:

Maurice Hilarius, Michal Jaegermann. Architecture:

Intel, Alpha and SPARC. Linux distribution:

Red Hat Linx and S.u.S.E. Special expertise:

Samba, Large capacity and high performance disk subsystems, RAID, data library applications, Digital Alpha systems, communications and networking application and setup.

Last modified:

March 6, 1997.

3.6 International Technology Sales

          500 North Cary Algonquin Road
          Cary, IL 60013
          U.S.A.

Phone:

+1 847 639 0775 Fax:

+1 847 639 0710 EMail:

rip@itsinfonet.com URL:

http://www.itsinfonet.com Contact:

Clarence Fleming Architecture:

Alpha. Linux distribution:

Red Hat Linux. Special expertise:

International Technology Sales is a Digital TOEM and Business Partner. We sell Digital Boards, Alpha CPU’s, Digital Computers and Systems and Digital Networking Products. As an authorized seller of Digital products we have been selling Boards and Systems for use with Linux.

Last modified:

March 6, 1997.

3.7 Linux Hardware Solutions, Inc.

     130 North Front Street
     Wilmington, NC 28401
     U.S.A.

Phone:

+1 888 LINUX HW, +1 910 251 9998 Fax:

+1 910 251 8832 EMail:

info@linux-hw.com URL:

http://www.linux-hw.com Contact:

Kit Cosper <cosper@linux-hw.com> Architecture:

Intel and Alpha. Linux distribution:

Red Hat Linux. Last modified:

March 6, 1997.

3.8 Nene Information Systems

          32 Glasgow Street
          St James
          Northampton
          NN5 5BL
          UK

Phone:

+44 1604 464561 Fax:

+44 1604 464561 EMail:

sales@nisys.demon.co.uk URL:

http://www.nisys.demon.co.uk Contact:

Timothy Peters <tim@nisys.demon.co.uk> Architecture:

Intel and Alpha. Linux distribution:

Linux Pro, Red Hat Linux and Slackware. Special expertise:

UK reseller for the Linux Mall and Cyclades Multiport Serial cards. Last modified:

March 10, 1997.

3.9 Netactive Systems, Ltd.

          19936 Roscommon Road
          Harper Woods, MI 48225-2254
          U.S.A.

Phone:

+1 810 582 0268 Fax:

+1 810 582 0268 EMail:

info@netactive.com URL:

http://www.netactive.com Architecture:

Intel, Alpha and SPARC. Linux distribution:

Every distribution available. Special expertise:

Sendmail, DNS, multihoming web and FTP servers: truly configured servers ready for your network. Last modified:

March 26, 1997.

3.10 Net Express

          3205 Isherwood Way
          Fremont, CA 94536
          U.S.A.

Phone:

+1 510 494 9493 (University or Government use only) Fax:

+1 510 790 2342 EMail:

sales@netex.tdl.com, tech@netex.tdl.com URL:

http://www.tdl.com/˜ netex Contact:

Roland Baker Architecture:

Intel and Alpha. Linux distribution:

Every distribution available. Special expertise:

High performance Linux workstations and SMP severs for U.S. Universities, U.S. Government agencies, scientists, engineers and the telecommunications industry. Rack mount chassis and pre-configured Internet servers and WAN hardware available. Our web site is a clearing house of technical information on designing and building high performance Linux systems from commodity PC hardware. Please see our web site for more information.

Last modified:

April 8, 1997.

3.11 Pacific Internet

     600 Corporate Pointe, Suite 100
     Culver City, CA 90230
     U.S.A.

Phone:

+1 310 410 9700 Fax:

+1 310 215 8071 EMail:

sales@pacnet.com URL:

http://www.pacnet.com Contact:

Tania Hayek, Michael Boyer, Denise Schutte, Maria Lepisto, Natalie Jackson Architecture:

Intel Linux distribution:

Caldera OpenLinux and Red Hat Linux.

Special expertise:

Internet Service Provider (ISP) systems. Last modified:

March 7, 1997.

3.12 PRONETICS

          M\"{a}rkische Stra{\ss}e 117
          D-44141 Dortmund
          Germany

Phone:

+49 231 557203 0 Fax:

+49 231 557203 27 EMail:

sales@pronetics.de URL:

http://www.pronetics.de Contact:

Horst Hohmeier Architecture:

Intel and PowerPC. Linux distribution:

GNU. Special expertise:

Internet server, proxy server, mail and news server, firewalls. Last modified:

April 11, 1997.

3.13 Sound Software Ltd.

          20 Abelard Avenue
          Brampton, ON L6Y 2K8
          Canada

Phone:

+1 905 452 0504 Fax:

+1 905 452 9754 EMail:

sales@telly.org URL:

http://www.telly.org/sound Contact:

Evan Leibovitch Architecture:

Intel. Linux distribution:

Caldera OpenLinux. Special expertise:

EDI, Progress DB, Internet integration, SCO/Unix migration. Authorized Caldera Business Partner. Last modified:

March 6, 1997.

3.14 SW Technology

          1701 North Greenville Suite 1104
          Richardson, TX 75081
          U.S.A.

Phone:

+1 972 907 0871 Fax:

+1 972 907 9339 EMail:

sales@swt.com URL:

http://www.swt.com Contact:

Marvin Y. Wu Architecture:

Intel, Alpha and SPARC. Linux distribution:

Caldera OpenLinux, Debian GNU/Linux and Red Hat Linux. Special expertise:

SW Technology is a fast growing company poised with products, services and expertise to address requirements for high performance computing, LAN and WAN access, dedicated synchronous connections, dial-in, dial-out, analog, digital and ISDN connections, and quality Internet services.

Last modified:

March 6, 1997.

3.15 Telenet System Solutions, Inc.

     2480 Kruse Drive
     San Jose, CA 95131
     U.S.A.

Phone:

+1 408 383 0334 x102 Fax:

+1 408 383 0335 EMail:

linux@tesys.com URL:

http://www.tesys.com Contact:

Billy Bath <billy@tesys.com> Architecture:

Intel and Alpha. Linux distribution:

Every distribution available.

Special expertise:

All of our systems are pretested and ready to go for the Internet. Our main client base is ISP’s and devlopers who want high end and reliable hardware.

Last modified:

March 26, 1997.